An online retailer of knee braces sells 22000 per month at a price of $60. They raise the price to $75 and sell only 16500 the next month.
a. Estimate the elasticity of demand for this price change, using the original P and Q as the reference.
b. What happened to total revenues as a result of this change?
c. Should the firm increase the price again, reverse the price increase, or leave the new price in place? Very brieflu explain.
Answer #1

A. Original price= $60

Original quantity= 22000

New price = $75

New Quantity= 16500

Elasticity of demand= (16500-22000)/(16500+22000)/2)÷(75-60)/(75+60)/2

=(--5500/19250)÷(15/67.5)

=-1.28

Demand is elastic.

b. Total revenue=Price* Quantity

Initial total revenue= 60*22000= $1320000

New total revenue= 75*16500=$1237500

As a result of the change, Total revenue fall.

c. Firm should reverse the price increase.

When demand is elastic, reducing the price leads to rise in total revenue. As the rise in quantity demanded is greater than the fall in price level.